OpenStudy (anonymous):

The plans of a public park include a circular fountain. The outline of the fountain can be modeled with the equation (x-4)^2 + (y+5)^2 = 50, where the units are meters and the graph is on a grid map of the park. a. Find the coordinates of the center. b. Find the circumference of the fountain. any help would be gladly appreciated xxx

OpenStudy (anonymous):

the circle equation\[\left( x - a \right)^{2} + \left( y - b \right)^{2} = r^{2}\]has center (a,b) and radius r

OpenStudy (anonymous):

Thank you :)